Three years ago Malcolm borrowed R7500 from Sarah on condition that he would repay her back in five years time. Interest is calculated at 13,5% every Three months.Nine months ago he also borrowed R2500 from her at an interest of 15,7% per year compounded monthly payable in two years from now. What will be the total amount that Malcolm owes Sarah two years from now?
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

looks like two separate accounts need to be set up.
one of them is at 13.5% per year compounded quarterly (every 3 months).
the other is at 17.5% per year compounded monthly.
for the first account, 7500 was borrowed 3 years ago and is due 2 years from now.
that's a total of 5 years * 4 = 20 quarters.
the formula to use is f = p * (1+r) ^ n
for the first loan, p = 7500, (1+r) = (1+.135/4) = 1.03375, n = 5 * 4 = 20
that's an interest rate of .03375 per quarter for 20 quarters.
the formula becomes f = 7500 * 1.03375 ^ 20 = 14567.05348.
for the second loan, p = 2500, (1+r) = (1+.157/12) = 1.013083333, n = 9 + 2*12 = 33.
that's an interest rate of .013083333 per month for 33 months.
the formula becomes f = 2500 * 1.013083333 ^ 33 = 3839.110744.
the total owed 2 years from now is 14567.05348 + 3839.110744 = 18406.16422.
round to the nearest penny to get 18406.16.
your monthly timeline is as follows:
a loan of 7500 3 years ago.
a loan of 2500 9 months ago.
a total of 18406.16 due 2 years from now.
let me know if you have any questions.
theo

Answer by ikleyn(52781)